# cp3-bench: A tool for benchmarking symbolic regression algorithms tested with cosmology

### v1

@@ 3 Benchmarking with cosmological datasets
Genetic Engine [GeneticEngine] GP No 2022
@@
PySR [PySR] GP,NN Yes 2023
@@
Table 2: Algorithms considered in this paper including the ML methods they use, if they support multi-core in the current implementation and the year of release. We denote Physics Inspired as PI, Neural Network as NN, Recurrent Neural Network as RNN, Reinforcement Learning as RL, and Genetric Programming as GP. AI-Feynman predominantly runs single-core. DSO and uDSR are not using GP methods due to a limitation of the current implementation.
@@
We bench the symbolic regression algorithms/models listed in Table 2. We assembled this collection of algorithms based on the availability of algorithms and a wish to consider algorithms based on a variety of methods. Additionally, we consider only algorithms that are generic in terms of the dataset types and thus do not consider algorithms which are specifically designed for, say, regression on redshift-distance data. Each model (except FFX) has several hyperparameters and other features which can be tuned. As described below, we have tuned the hyperparameters using a set of seven standard datasets with one to three features each. These datasets were chosen among standard datasets used for tuning hyperparameters of symbolic regression algorithms and are listed in Appendix LABEL:app:StandardDataset [makke2023interpretable]. Tuning hyperparameters on a generic test dataset is a simple way of avoiding biasing for/against any of the algorithms or datasets in the tuning. We simply tune using standard datasets used in the symbolic regression community and which have no relation to cosmological datasets (see Subsection 3.1.3 for more details on the hyperparameter tuning). The tuned hyperparameters can be found in the publicly available code where all the algorithms are in the methods folder and hyperparameters are set in procedure.py (see Figure 1 for details). Note that the algorithms DSO and uDSR in principle support Genetic Programming (GP) methods but the interface we use currently using does not support this feature.
@@
Regarding the algorithms we note that DSR and uDSR are further developments of DSO. From the test dataset we find that these algorithms are in some cases indeed able to find the correct expression. FFX is fast, has no hyperparameters and would only be able to provide the correct symbolic equation in very limited cases. GPZGD and Genetic Engine were hard to tune and the former did not recover any of our test equations. AI-Feynman is rather slow, but was able to find one equation. Of the purely genetic algorithms QLattice and ITEA perform the best. PySR is also one of the more successful algorithms according to the training done on the test datasets.
@@
The authors of this paper are aware of the existence of the benchmark tool SRBench [lacava2021contemporary] and have used it as inspiration. Due to e.g. technical issues, we found it easier to make a new tool, and create our own documentation. The main guide for designing the tool has been to make it as user-friendly as possible both in terms of implementation and adding extensions.
@@ 3.1 Benchmarking method
In this section we discuss the benchmark tools developed for this paper and the procedure we have used to achieve the results presented in Section 3.2.
@@ 3.1.1 Framework
As part of this paper we release the code cp3-bench and Things-to-bench used for the benchmarking presented in this paper. The release contains two components. Firstly, we are releasing the benchmark engine which we call cosmology and particle physics phenomena benchmark (cp3-bench). Secondly, we release our Things-to-bench wrapper repository showcasing how one can leverage the benchmark tool in an easy way, and create a simple script to loop over datasets of interest. The latter repository contains the datasets used here, the code used to run them, as well as the raw results including mathematica notebooks used to simplify and prepare the data for presentation in Appendix LABEL:app:full-results. The code can be found at:
Datasetsmain.pyThings-to-benchevaluate_dataset.pyresults/result.csvBenchmarkalgorithmsinstall.pycp3-benchprocedure.py__init__.pyrequirements.txtconfig.jsontestsAlgorithm folder/benchCLICLIpyenv virtualenv
@@
We release cp3-bench with ten algorithms implemented and readily available for use. Adding new algorithms is straightforward by creating a new folder in the methods folder with the structure described in Figure 1. The program uses the folders in methods to recognize which methods are implemented. The default is to install all models when running the module install.py. Both modules can be imported to other scripts or run with CLI (command-line interface) and has the optional flag ---methods which allows the user to select specific algorithms to install or benchmark. The install module also allows for the optional flag ---reinstall which by default is false, but can be chosen to be true to reinstall a misbehaving algorithm. The module evaluate_dataset.py requires an argument specifying the path to a .csv dataset to evaluate. This flag is called ---data_path.
@@
The tests folder contains the configuration of universal tests, which each algorithm should pass. During installation these are used to verify the functioning of the algorithms.
@@
The most important file is the procedure.py script. Here the algorithm is imported and defined according to the framework. It is also where the hyperparameters are set. The script defines a class which initializes the algorithm and defines a procedure for training this algorithm. Note that some algorithms have a format_output function to reformat the output equation. Some implementations may also drop coefficients that are small or set small numbers to zero and round numbers to fewer significant digits to produce a shorter equation that is easier to read. In some cases, this formatting turns the symbolic expression found by the algorithm into the ground truth. In such a case, the MSE reported will not be zero, even though the output is exactly correct since the MSE is based on the unformatted result which contained small (errornous) additional terms. Some of the methods cp3-bench import have similar features built into the algorithm, one example is uDSR which can drop coefficients smaller than a specified value.
The algorithm classes inherit elements from a global superclass called MethodEvaluator. Most importantly, this class manages some logging levels and the evaluate method that is run during benchmarks. Specifically, it is run by the CLI interface that is used by cp3-bench to interact between the algorithms and the evaluate_dataset.py module.
@@
3.1.2 Getting started
@@
Since this package relies on pyenv to manage virtual environments, it is key that this dependency is correctly installed and initialized. Otherwise you can face a wide variety of issues as mentioned in the README.md file on the GibHub repository page.
@@
The result of the installation will be logged in the STATUS.md file in the bench folder. You can verify that everything is installed correctly by checking this file. There are four columns to consider:
  • Config: Checks the configuration is valid
  • Environment: Checks the environment is correctly setup
  • Tests: Checks that all tests pass successfully
To use the package you need to parse a .csv file with the output value column named “target”. You can find short example of how this looks in the tests/utils folder by looking at the test_dataset.csv file.
@@
1from python.bench.evaluate_dataset import evaluate_dataset
@@
and then parse the input to it as a regular Python module in/out script in Things-to bench. Alternatively you can also use CLI to evaluate a dataset using the following command from the root folder of cp3-bench:
@@
Another option is to specify which algorithms to use by setting the appropriate flag, as with the installation command. The output is in the results folder.
@@ 3.1.3 Benchmark procedure
The primary goal of this paper is to evaluate the performance of symbolic regression models on cosmological datasets. Many of the models we include in the benchmark were originally tested by their creators on artificial standard datasets without direct application to cosmology or astrophysics, often making the regression task much easier. Therefore, we begin by considering a small sample of standard datasets, which can be found in Appendix LABEL:app:StandardDataset.
We optimize the hyperparamters to perform well across these standard datasets. Since the algorithms use different ML methods, it is not straightforward to compare the models one-to-one. For instance, certain models may be slower than others, while some models do not benefit from extra time, as they have reached a minimum. It is therefore not generally meaningful to aim for a similar runtime. Runtime is nonetheless one quantitative measure of the usefulness of the model and we will therefore include it when comparing the algorithms below.
@@
where $n$ is the number of values in the output vector $y$ and the prediction vector $y_{i}$.
We can thereby recognize that an algorithm may be somewhat slow compared to others, but if the symbolic expression is correct, then this is generally more important than having the model completing faster. On the other hand, if two models are equally good or poor at obtaining the correct symbolic expression, then a comparison of run times becomes relevant for differentiating the algorithms. The MSE is simply a measure of the global fit of the symbolic expressions. If an algorithm obtains the correct symbolic equation, the MSE will vanish. For all other situations, the MSE can be considered a measure of how far the resultant symbolic expression is from producing a meaningful result and possible ground truth. But one should remember that a low MSE is not necessarily equal to being close to finding the ground truth since an inherently incorrect expression may be fine-tuned to fit the data well within the training region.
@@
      The resultant equation is equal to the expected, possibly written in a different way. Consider the example where the expected equation is,
      $f(x)=x^{2}+1$ (3.2)
      but the algorithm resulted in,
@@
      The coefficients of the result are correct to the first significant digits. Assuming the equation from above, we would accept the following as correct,
      $f(x)=\frac{1.1x^{3}+x}{x}.$ (3.4)
@@
      $f(x)=x^{2}-10^{-10}x+1\rightarrow f(x)=x^{2}+1,$ (3.5)
      is considered correct.
  • ✓(MSE): We give a check-mark but with the MSE in parenthesis to indicate that the algorithms derived the correct equation at least structurally, but there are larger deviations than for the correct case such as:
@@
      The coefficients are not correct to the first significant digit. In this category, the result must still contain all the correct terms but there may be an additional term with a significantly smaller coefficient in front. Continuing with Equation 3.2 as the correct function, we would in this category allow larger deviations of the coefficient:
@@
      Residual terms that are less significant than the main terms of the result are accepted. For instance,
      $f(x)=x^{2}-0.0001x+1,$ (3.7)
@@
      $f(x)=x^{2}-x+1,$ (3.8)
      is, however, not accepted in this category since the coefficient of the extra term is of the same order as the correct terms and the extra term is no longer considered a residue.
  • MSE: Any expression deviating from the ground truth in ways not described above are considered incorrect and we indicate the deviation in terms of the MSE.
After tuning the models on the standard dataset (the result can be seen in Table LABEL:fig:test-results), we run the cosmological datasets on the models and evaluate the results. This means our fine-tuning is blind to the cosmological dataset and we have only seen the results from benchmarking the cosmological data after we have locked the hyperparameters. In this way one can test the meaningfulness of optimizing for standard datasets and how well the models perform on cosmological datasets. The hardware used is described in Table 3.
@@
Out of the box, the method evaluate_dataset.py will run multiple algorithms parallel, which might lead to some fight over resources if CPU is overloaded, but overall it will speed up the procedure and yield results faster. For the testing we kept the code as it is out of the box, but for the benchmarking of the cosmological dataset we specifically run only one algorithm per dataset at a time. This can be done by modifying the code, or by simply selecting to only use one algorithm at a time, instead of the default which is to use all algorithms.

### v2

@@ 3 Benchmarking with cosmological datasets
In this section we provide the details of cp3-bench and our demonstration benchmark. We begin by giving an overview of the 12 algorithms included in the benchmark and discussing their hyperparameter tuning.
@@
Genetic Engine (GE) [GeneticEngine] GP No 2022
GPG [gpg] GP Yes 2022
@@
Operon [operon] GP Yes 2020
PySR [cranmer2023interpretable] GP,NN Yes 2023
@@
Table 2: Algorithms considered in this paper including the ML methods they use, if they support multi-core in the current implementation and the year of their release. We denote Physics Inspired as PI, Neural Network as NN, Recurrent Neural Network as RNN, Reinforcement Learning as RL, and Genetric Programming as GP. AI-Feynman predominantly runs single-core. DSO and uDSR are not using GP methods due to a limitation of the current implementation of cp3-bench (cp3-bench uses the scikit-learn Python wrapper but as stated in the documentation for DSO and uDSR, GP is not supported when using this wrapper).
@@
We bench the symbolic regression algorithms/models listed in Table 2. We assembled this collection of algorithms based on the availability of algorithms and a wish to consider algorithms based on a variety of methods. Each model (except FFX) has several hyperparameters and other features which can be tuned. As described below, we have tuned the hyperparameters using a set of seven standard datasets with one to three features each. These datasets were chosen from [makke2023interpretable] among standard datasets used for tuning hyperparameters of symbolic regression algorithms and are listed in Appendix LABEL:app:StandardDataset. Tuning hyperparameters on a generic test dataset rather than the cosmological datasets described above is a simple way of avoiding biasing for/against any of the algorithms or datasets. We simply tune using standard datasets used in the symbolic regression community and which have no relation to cosmological datasets (see Section 3.1.4 for more details on the hyperparameter tuning). The tuned hyperparameters can be found in the publicly available code where all the algorithms are in the methods folder and hyperparameters are set in procedure.py (see Figure 8 for details). Note that the algorithms DSO and uDSR in principle support Genetic Programming (GP) methods but the interface we use currently does not support this feature.
@@
Regarding the algorithms we note that DSR and uDSR are further developments of DSO. From the test dataset we find that these algorithms are in some cases indeed able to find the correct symbolic expressions corresponding to the datasets. FFX is fast, has no hyperparameters and would only be able to provide the correct symbolic equation in very limited cases where only multiplication, polynomials, and division is used. GPZGD and Genetic Engine (GE) were hard to tune and the former did not recover any of our test equations. GPG is reasonably fast but struggles with rather simple equations. AI-Feynman is rather slow, but was able to find one correct symbolic expression for the test data. Of the purely genetic algorithms Operon, QLattice, and ITEA perform the best on the test data. PySR is also one of the more successful algorithms according to the training done on the test datasets.
In the following subsections we detail the procedure of benchmarking with cp3-bench and Things-to-bench. We first introduce the framework for cp3-bench and installation options before moving on to discussing the specific procedure (rules and criteria) we applied in our demonstration benchmark.
@@
We are aware of the existence of the benchmark tool SRBench [lacava2021contemporary] and have used it as inspiration. The main guide for designing our tool has been to make it as user-friendly as possible both in terms of implementation and adding extensions. Since SRBench has come to represent the standard with which SR algorithms are benched by developers, we compare cp3-bench and SRBench at the end of the section.
@@ 3.1 Benchmarking method
In this subsection we discuss the benchmark tools developed for this paper and the procedure we have used to achieve the results presented in Section 3.2.
@@ 3.1.1 Framework
As part of this paper we release the code cp3-bench and Things-to-bench used for the benchmarking presented in this paper. The release contains two components. Firstly, we are releasing the benchmark engine which we call cosmology and particle physics phenomenology benchmark (cp3-bench). Secondly, we release our Things-to-bench wrapper repository showcasing how one can leverage the benchmark tool in an easy way, and create a simple script to loop over datasets of interest. A user can also use cp3-bench as project folder and put in the datasets and utility scripts straight into the cp3-bench folder and build a Docker image on that. The disadvantage of this is that changing versions of cp3-bench is less convenient as cp3-bench would not be separated from the datasets and other scripts in this case. The latter repository contains the datasets used for this work, the code used to manage the benchmark of them, as well as the raw results including Mathematica notebooks used to simplify and prepare the data for presentation in Appendix LABEL:app:full-results. The code can be found at:
Datasetsmain.pyThings-to-benchevaluate_dataset.pyresults/result.csvmethodsinstall.pycp3-benchprocedure.py__init__.pyrequirements.txtconfig.jsontestsAlgorithm folder/benchCLICLIpyenv virtualenv
@@
We release cp3-bench with 12 algorithms implemented and readily available for use. Adding new algorithms is straightforward by creating a new folder in the methods folder with the structure described in Figure 8. The program uses the folders in methods to recognize which methods are implemented. The default is to install all models when running the module install.py. Both modules can be used by other scripts or run with CLI (command-line interface) and have the optional flag ---methods which allows the user to select specific algorithms to install or benchmark. The install module also allows for the optional flag ---reinstall which by default is false, but can be chosen to be true to reinstall a misbehaving algorithm. The module evaluate_dataset.py requires an argument specifying the path to a .csv dataset to evaluate. This flag is called ---data_path.
@@
The tests folder contains the configuration of universal tests, which each algorithm should pass. During installation these are used to verify the functioning of the algorithms. A summery of the installation is logged in a file called STATUS.md in the bench folder which indicate if there were any failures during the installation. You can verify that everything is installed correctly by checking this file. There are four columns to consider:
  • Config: Checks the configuration is valid
  • Environment: Checks the environment is correctly setup
  • Tests: Checks that all tests pass successfully
@@
The most important file is the procedure.py script. Here the algorithm is imported and defined according to the framework. It is also where the hyperparameters are set. The script defines a class which initializes the algorithm and defines a procedure for training this algorithm. Note that some algorithms have a format_output function to reformat the output equation. Some implementations may also drop coefficients that are small or set small numbers to zero and round numbers to fewer significant digits to produce a shorter equation that is easier to read. In some cases, this formatting turns the symbolic expression found by the algorithm into the ground truth. In such a case, the MSE reported will not be zero, even though the output is exactly correct since the MSE is based on the unformatted result which contained small (errornous) additional terms or deviations to coefficients. Some of the methods cp3-bench imports have similar features built into the algorithm, one example being uDSR which can drop coefficients smaller than a specified value.
The algorithm classes inherit elements from a global superclass called MethodEvaluator. Most importantly, this class manages some logging levels and the evaluate method that is run during benchmarks. Specifically, it is run using Command Line Interface (CLI) that is used by cp3-bench to interact between the algorithms and the evaluate_dataset.py module.
@@
3.1.2 Getting started with a manual installation
@@
Since this package relies on pyenv to manage virtual environments, it is key that this dependency is correctly installed and initialized. Otherwise you can face a wide variety of issues as mentioned in the README.md file on the GibHub repository page. Start by cloning the repository:
1git clone https://github.com/CP3-Origins/cp3-bench.git
@@
To use the package you need to parse a .csv file with the output value column named “target”. You can find a short example of what this looks like in the tests/utils folder by looking at the test_dataset.csv file.
@@
1from cp3bench.bench.evaluate_dataset import evaluate_dataset
@@
and then parse the input to it as a regular python module in our script in Things-to bench. Alternatively you can use the CLI to evaluate a dataset using the following command from the root folder of cp3-bench:
@@
Another option is to specify which algorithms to use by setting the appropriate flag, as with the installation command. The output is in the results folder. For research projects we recommend using cp3-bench as an imported submodule to another repository managing the benchmark project as we have done in Things-to-bench.
3.1.3 Getting started with an automated installation
To give the user an out-of-the-box experience we also offer a Dockerfile to support an automated installation with Docker. One disadvantage with using a Dockerfile is that you need to compile a new docker image to update your project, and making the same integration as we have done with Things-to-bench will require some reconfiguration by the user. Users who are simply interested in leveraging cp3-bench for benchmarking can simply clone the project, add a new folder with the datasets of interest (and potentially a script to manage them) before creating the Docker image. The added datasets will then be available to benchmark in the Docker image. To install the benchmark in an automated fashion you need:
  • Docker
Start by cloning the repository:
1git clone https://github.com/CP3-Origins/cp3-bench.git
For compatibility reasons we cannot used Docker BuildKit, which is enabled per default on the latest version. Ensure BuildKit is disabled by setting the following environmental variable: DOCKER_BUILDKIT=0. With this done and any dataset added to your cp3-bench folder you can build the package using:
1docker build . -t cp3-bench
As an alternative to setting DOCKER_BUILDKIT=0 as an environmental variable, you can also specify it during the build using the command:
1DOCKER_BUILDKIT=0 docker build . -t cp3-bench
If you are using MacOS with ARM chips you need to use x86 emulation. This can be done by adding the following parameter --platform=linux/amd64. In this case the command including disabling BuildKit becomes:
1DOCKER_BUILDKIT=0 docker build . --platform=linux/amd64 -t cp3-bench
Like with other commands of this tool, the default is to install all algorithms but you can specify only a subset of algorithms to install. For the Docker build you can use the following syntax to specify the subset with a comma-separated list:
1docker build . -t cp3-bench --build-arg METHODS=ffx,gpg
Here METHODS is the argument which stores what algorithms to install, the default is all, but in this example two algorithms are choosing namely ffx and gpg.
Depending on how many algorithms you install and the hardware you use, the installation may take above 30 minutes. Upon completion you can setup and enter a Docker container with your fully functional benchmark tool using the command:
1docker run -it cp3-bench
This will mount your terminal to that of the Docker container and you are now in the isolated environment of the cp3-bench package with any additional files you have added available to bench. For Mac on ARM chips this becomes:
1docker run --platform=linux/amd64 -it cp3-bench
Early chips like the M1 does not support DSO, DSR, Operon, PySR and uDSR. We expect the remaining methods would work on these chips.
@@ 3.1.4 Benchmark procedure
The primary goal of this paper is to evaluate the performance of symbolic regression models on cosmological datasets. Most of the models we include in the benchmark were originally tested by their creators on artificial standard datasets without direct application to cosmology or astro(particle) physics. It is not clear that performance on standard datasets can in general be expected to be representative of performance on real datasets e.g. representing cosmological data. We will therefore also compare the performance of the SR algorithms on the standard test datasets that we use for hyperparameter tuning with their performance on the cosmological datasets.
We optimize the hyperparamters to perform well across the standard test datasets listed in Appendix LABEL:app:StandardDataset. The metric for successful tuning has been to increase the number of successfully identified symbolic expressions. Since the algorithms use different ML methods, it is not straightforward to compare the models one-to-one. For instance, certain models may be slower than others, while some models do not benefit from extra time, as they have reached a minimum. It is therefore not generally meaningful to aim for a similar runtime. Runtime is nonetheless one quantitative measure of the usefulness of the model and we will therefore include it when comparing the algorithms below.
@@
where $n$ is the number of values in the target vector $y_{i}$, and $\hat{y_{i}}$ represents the corresponding SR algorithm prediction.
We can thereby recognize that an algorithm may be somewhat slow compared to others, but if the symbolic expression is correct, then this is generally more important than having the model completing faster. On the other hand, if two models are equally good or poor at obtaining the correct symbolic expression, then a comparison of run times becomes relevant for differentiating the algorithms. The MSE is simply a measure of the global fit of the symbolic expressions. If an algorithm obtains the correct symbolic equation and the data has no error, the MSE will vanish. For all other situations, the MSE can be considered a measure of how far the resultant symbolic expression is from producing a meaningful result and possible ground truth (for data without error). One should nonetheless remember that a low MSE is not necessarily equal to being close to finding the ground truth since an inherently incorrect expression may be fine-tuned to fit the data well within the training region. The MSE can additionally become misleading for datasets with errors since a ground truth will clearly not in general be achieved by having the lowest possible MSE. We also emphasize that the symbolic expressions identified by the algorithms are generally not obtained through only considering the MSE of the expressions. Several of the algorithms for instance also take the complexity of the expressions into account, preferring simpler expressions over more complex ones. We do not implement any selection strategy into cp3-bench but simply pull out the symbolic expression each algorithm has ranked as “the best”, using their individual built-in decision rules. We are thus simply including the MSE here as a simple way of comparing the obtained expressions.
@@
      The resultant equation is equal to the expected, possibly written in a different way. Consider the example where the expected equation is
      $f(x)=x^{2}+1.$ (3.2)
      Assume the algorithm resulted in
@@
      The coefficients of the result are correct to the first significant digits. Assuming the equation from above, we would accept
      $f(x)=\frac{1.1x^{3}+x}{x}$ (3.4)
      as correct.
@@
      $f(x)=x^{2}-10^{-10}x+1\rightarrow f(x)=x^{2}+1$ (3.5)
      is considered correct. In cases where the ground truth is unknown, a ✓will not be given unless the MSE is negligible both inside the training region and on any reasonable parameter region beyond the training region. The reasonable parameter region must be determined for each relevant dataset, based on what parameter regions are physically sensible. Only in such a case could we claim to believe we have identified a new physical relation.
  • ✓(MSE): We give a check-mark with the MSE in parenthesis to indicate that the algorithms derived the correct equation at least structurally, but there are larger deviations than for the correct case defined above, such as:
@@
      The coefficients are not correct to the first significant digit. In this category, the result must still contain all the correct terms but there may be an additional term with a significantly smaller coefficient in front. Continuing with Equation 3.2 as the correct function, we would in this category allow larger deviations of the coefficient such as
@@
      Residual terms that have smaller prefactors than the main terms of the result are accepted. For instance,
      $f(x)=x^{2}-0.0001x+1$ (3.7)
@@
      $f(x)=x^{2}-x+1$ (3.8)
      is, however, not accepted in this category since the coefficient of the extra term is of the same order as the correct terms and the extra term is no longer considered a residue. In cases where we do not know the ground truth, a “✓(MSE)” is given if the MSE is sub-percent both inside the training region and on a parameter region that extends to the main part of the physically reasonable parameter region. In this case we expect that the expression found may be close to representing a genuine physical relation, while it being clear that an exactly correct expression has not been identified.
  • MSE: Any expression deviating from the ground truth in ways not described above are considered incorrect and we indicate the deviation in terms of the MSE. For cases where the ground truth is not known, we present the MSE when the expression cannot be categorized in either of the above categories.
A scheme for categorizing expressions will inevitably contain some arbitrariness unless we very strictly consider an expression correct only if it is exactly so and otherwise consider the expression incorrect. However, such a very strict scheme seems unfair and prone to misleading categorizations and can easily become biased because some of the SR algorithms themselves contain e.g. round offs. Although there is thus some arbitrariness to our choices such as ignoring numbers smaller than $10^{-10}$, our opinion is that such a scheme is more appropriate than a more strict one. If one is trying to discover new equations with SR methods, the deviations in our “almost correct” category would presumably be detected manually if the general structure of the symbolic expression is correct. Note also that we show the symbolic expressions identified by each SR algorithm for each dataset in Appendix LABEL:app:full-results and the reader may consult that appendix to get a fuller impression of the applicability of each expression identified by the 12 algorithms for the 28 datasets.
In the results section we also include a check-mark in parenthesis, (✓), which is particularly relevant for the datasets in group $C5$. In this case the parenthesis indicates that the result contains a constant pseudo-parameter, that ideally shouldn’t have been in the expression, but evaluating this constant will yield the correct result. E.g. consider the superficial constant parameter $a=1$. In the case that the true expression is $f(x)=x^{2}+1$, but the output expression is $f(x)=x^{2}+a$, then setting $a\rightarrow 1$ clearly yields the correct equation, but we show that the algorithm failed to identify this pseudo-parameter as something to be ignored and replaced with a constant by marking it with (✓).
After tuning the models on the standard dataset (the result can be seen in Table LABEL:fig:test-results), we run the cosmological datasets on the models and evaluate the results. Note specifically that our hyperparameter fine-tuning is blind to the cosmological datasets and that we have only seen the results from benchmarking the cosmological data after we have locked the hyperparameters. This way one can test the meaningfulness of optimizing for standard datasets and how well the models perform on cosmological datasets. The hardware used is described in Table 3.
@@
Out-of-the-box, the method evaluate_dataset.py will run multiple algorithms in parallel, which might lead to some fight over resources if CPU is overloaded, but overall it will speed up the procedure and yield results faster. For the testing we kept the code as it is out-of-the-box, but for the benchmarking of the cosmological dataset we specifically run only one algorithm per dataset at a time. This can be done by modifying the code, or by simply selecting to only use one algorithm at a time, instead of the default which is to use all algorithms.
We note that initially cp3-bench (v1.0.0) only came with ten algorithms. Due to feedback from an anonymous referee and a code developer, we added two additional algorithms, Operon and GPG, as part of a revised version of this work. These two algorithms where implemented in the same fashion as the original ten, where we first optimize on the test datasets and then, after fine-tuning, apply them on the cosmological datasets. However, when benchmarking the first ten algorithms we gained experience of how models optimized on the training set behave on the cosmological datasets. For instance, we became aware of how strong polynomial performance is important for performing well on the cosmological datasets. This knowledge may subconsiously have affected the hyperparameter tuning of the two newly added algorithms Operon and GPG. Since we have also added nine additional datasets to the benchmark, we expect that at least for these datasets this potential bias has a vanishing effect.
3.1.5 Comparison with SRBench
As stated on the homepage of SRBench [^8] https://cavalab.org/srbench/ [lacava2021contemporary], the main motivation for that tool is to make a unified platform and framework for SR and genetic programming. This includes having a single platform with a large dataset to perform uniform benchmarking on the considered algorithms. SRBench therefore largely serves the needs of SR developers and thus what we consider more “academic” studies into SR algorithms. Our tool, cp3-bench, aims at easing the practical use of multiple SR algorithms simultaneously and especially in making many algorithms easily accessible to researchers not actively developing SR algorithms but interested in their application to their given research tasks. We believe one big obstacle in choosing the most optimal SR algorithm for this group of researchers is installing the algorithms. Most of the algorithms we considered (but not all) are in themselves fairly difficult to install. This in itself is a hindrance for researchers to consider a broad range of SR algorithms. The hindrance is significantly exacerbated by various algorithms requiring different versions of e.g. Python and TensorFlow. For instance some algorithm require older Tensorflow 1.x versions which are incompatible with Tensorflow 2.x versions. To improve user experience, we handle the setup of individual environments for each algorithms for the user.
While SRBench provides environments for the individual algorithms to handle dependencies, there may still be some difficulties in archiving a successful installation, and there is no easy way to install all algorithms in one go and benchmark a dataset on several methods using SRBench. In addition, older algorithms such as FFX currently do not work on SRBench due to a deprecation issue with the FFX package. With cp3-bench we have ways to patch such issues and can thus present a working version of the FFX algorithm.
To complement the SRBench setup and make SR algorithms more easily available to potential users of SR algorithms, we provide cp3-bench in a format where the user simply installs pyenv  [^9] https://github.com/pyenv/pyenv and afterwards the installation of cp3-bench automatically includes the installation of all algorithms unless the user actively chooses only to install some (or none) of the algorithms. Thus, a main difference between SRBench and cp3-bench is that the former is aimed at supplying a benchmark while the latter is aimed at being a tool that makes using and comparing SR algorithms easy for non-developers. If the user has Docker it is possible to install everything with all dependencies robustly set with a single command to provide a true out-of-the-box user experience. Since we are not specifically aiming cp3-bench at the SR developer community, we decided to use a data format we expect the cosmology and astroparticle physics community will be well-familiar with and find easy and intuitive to use. This is also why we make use of simple csv files as input and output, but there is no reason why one could not develop PMLB [^10] PMLB (Penn Machine Learning Benchmark, https://epistasislab.github.io/pmlb/, is a large ML database with a different file format than that used here. support if one wishes to leverage this.